# Anomalous and linear holographic hard wall models for light unflavored mesons

### v1

@@ III Anomalous HW model
In an interacting theory, the dimension of an operator is no longer canonical, thanks to renormalization that implies in anomalous contributions ($\Delta_{\rm Anom}$). So, the total dimension of an operator is given by its canonical plus the anomalous dimensions:
@@
From a well known semi-classical gauge/string duality analysis Gubser:2002tv , it is found that the anomalous dimension increases logarithmically with the spín $S$ of the boundary operators as
@@
where $\lambda$ is ’t Hooft coupling. Inspired by this result, we consider an anomalous dimension given by
$\Delta_{\rm Log}=a\ln(S),$ (22)
where $a$ is a constant that will be determined by best fit for each meson family.
@@
$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm QCD};\qquad\qquad(\nu=S+1+\Delta_{\rm Log}),$ (23)
where we used the fact that $\Delta_{\rm can.}=3+S$ for these mesonic states and that the order of the corresponding Bessel function is $\nu=\Delta_{\rm Total}-2$. The value of $\Lambda_{\rm QCD}$ is determined for each meson familiy using as inputs the masses of the states with $S=1$ from PDG, namely $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$. These values of $\Lambda_{\rm QCD}$ coincide with the ones used in the usual HW model, listed in Table 2, once Eq. (22) implies a vanishing anomalous dimension for $S=1$.
In order to obtain the best numerical fit, let us define the percentual deviation:
$q_{i}=\frac{|m_{i}-M_{i}|}{M_{i}},$ (24)
@@
$Q=\sqrt{\sum_{i}q_{i}^{2}}.$ (25)
So, the best fit method adopted here corresponds to minimize the function $Q$ for each meson family in the AHW model.
We will also consider an uncertainty for the parameter $a$, defined in Eq. (22), which implies in uncertainties for the masses ${m_{i}}$, obtained by error propagation, accordingly to the formula:
@@
where $\delta m_{0}$ and $\delta\chi_{\nu_{i},1}$ are the uncertainties in the mass of the state with $S=1$ and in the zero of Bessel function that is related to the mass $m_{i}$. The uncertainty $\delta m_{0}$ of the state with $S=1$ for each meson family in our model will be the arithmetic average of the percentual uncertainties of PDG masses $M_{i}$ for that given family. Defined our criterion for best fit and how we compute the uncertainties $\delta{m_{i}}$ for each mass $m_{i}$, we are able to tackle with the AHW model for light unflavored mesons.
III.1 AHW model for mesons with even ${\cal CP}$, $I=1$, and $S\geq 1$
We start applying the AHW model for mesons with even ${\cal CP}$, isospin $I=1$, and spins $S\geq 1$, taking the mass of the $\rho(770)$ state from PDG as an input, which fixes $\Lambda_{\rm QCD}=150.96$ MeV. We minimize the parameter $Q$, given by Eq. (25), with respect to PDG masses from Table 1, finding the logarithm anomalous dimensions, Eq. (22), to be
@@ III.1 AHW model for mesons with even ${\cal CP}$, $I=1$, and $S\geq 1$
The masses with its uncertainties obtained within this model, from Eq. (23), with these anomalous dimensions, are presented in Table 3, together with percentual deviations in comparison with PDG data. The value of $\Delta_{\rm Log}(S)$ of the above equation for each spin $S$ is also shown in the last column of this Table. The Regge trajectory obtained by doing linear regression with the obtained masses from this AHW Log model is given by
$J=(0.88\pm 0.01)M^{2}+0.50\pm 0.03,$ (28)
@@
with $\chi^{2}/ndf=4.32/4=1.08$. This trajectory is presented in the upper left panel of Fig. 1 as the blue line, and can be compared with the Regge trajectory of this family from PDG data, shown in Table 2. The red dots are the values in the $J\times M^{2}$ plane of the mesons $\rho(770)$, $a_{2}(1320)$, $\rho_{3}(1690)$, $a_{4}(1970)$, $\rho_{5}(2350)$, and $a_{6}(2450)$ in this AHW Log model, with their uncertainties.
III.2 AHW model for mesons with even ${\cal CP}$, $I=0$, and $S\geq 1$
Here, we discuss the case of mesons with even ${\cal CP}$, isospin $I=0$, and spins $S\geq 1$. Minimizing $Q$, given by Eq. (25), we obtain $a=2.61\pm 0.09$ for the anomalous dimension, Eq. (22), which reads
$\Delta_{\rm log}(S)=(2.61\pm 0.09)\ln(S).$ (29)
Using the mass of the state $\omega(782)$ from PDG as an input, which implies $\Lambda_{\rm QCD}=152.24$ MeV for this case, we calculate the tower of masses with the uncertainties in this AHW Log model and show these results in Table 3. The percentual deviations of these masses in comparison with PDG are also presented, as well as the value of the anomalous dimension $\Delta_{\rm Log}$, for each meson state. The Regge trajectory obtained by doing linear regression with the masses obtained by the AHW Log model is given by
$J=(0.90\pm 0.01)M^{2}+0.48\pm 0.03,$ (30)
with $\chi^{2}/ndf=8.88/4=2.22$. This trajectory is presented in the upper right panel of Fig. 1 as the blue line. The red dots are the values in the $J\times M^{2}$ plane of the mesons with even ${\cal CP}$, isospin $I=0$, and spins $S\geq 1$, in this AHW Log model, with their uncertainties. This trajectory can be compared with the one obtained from PDG for the $\omega(782)$ family, presented in Table 2.
III.3 AHW model for mesons with odd ${\cal CP}$, $I=1$, and $S\geq 1$
Now we present our results for mesons with odd ${\cal CP}$, $I=1$, and $S\geq 1$. The minimization of $Q$, Eq. (25), gives $a=0.56\pm 0.03$ for the the anomalous dimension, Eq. (22), which becomes
$\Delta_{\rm Log}(S)=(0.56\pm 0.03)\ln(S).$ (31)
The masses obtained in this AHW Log model are shown in Table 3, with their uncertainties, where we used the mass of the state $b_{1}(1235)$ as an input, so that $\Lambda_{\rm QCD}=239.41$ MeV. The Regge trajectory obtained by linear regression with these masses is given by
$J=(0.78\pm 0.02)M^{2}-0.13\pm 0.08,$ (32)
with $\chi^{2}$/ndf$=27/2=13.5$. This trajectory is represented in the lower left panel of Fig. 1 as the blue line and can be compared with the one obtained from PDG data, shown in Table 2. The red dots are the values obtained from this AHW Log model, with their corresponding uncertainties.
Table 3: Masses in MeV for mesons classified under even or odd $\mathcal{CP}$ with isospin $I=0,1$ and spins $S\geq 1$ from the AHW Log model, determined from Eq. (23), and the percentual deviations $\delta_{\rm Log}$ of the masses from this model compared with PDG data. For each family the mass of the state with $J=S=1$ is taken as an input from PDG. In the last column we show the anomalous dimension $\Delta_{\rm Log}$ for each state determined by Eq. (22).
$J^{PC}$ $\mathcal{CP}$ $I$ Meson AHW Log $\delta_{\rm Log}$ $\Delta_{\rm Log}$
$1^{--}$ even 1 $\rho(770)$ $775.26\pm 10.08$ $0\%$ 00
$2^{++}$ even 1 $a_{2}(1320)$ $1309.95\pm 22$ $0.53\%$ $1.92$
$3^{--}$ even 1 $\rho_{3}(1690)$ $1681.02\pm 31$ $0.46\%$ $3.04$
$4^{++}$ even 1 $a_{4}(1970)$ $1988.75\pm 37$ $1.11\%$ $3.84$
$5^{--}$ even 1 $\rho_{5}(2350)$ $2262.37\pm 43$ $2.90\%$ $4.46$
$6^{++}$ even 1 $a_{6}(2450)$ $2514.58\pm 47$ $2.64\%$ $4.96$
$1^{--}$ even 0 $\omega(782)$ $782.66\pm 7.83$ $0\%$ 00
$2^{++}$ even 0 $f_{2}(1270)$ $1302.58\pm 17$ $2.12\%$ $1.81$
$3^{--}$ even 0 $\omega_{3}(1670)$ $1666.4\pm 23$ $0.04\%$ $2.87$
$4^{++}$ even 0 $f_{4}(2050)$ $1969.63\pm 28$ $2.40\%$ $3.62$
$5^{--}$ even 0 $\omega_{5}(2250)$ $2240.19\pm 32$ $0.44\%$ $4.20$
$6^{++}$ even 0 $f_{6}(2510)$ $2490.23\pm 36$ $1.02\%$ $4.68$
$1^{+-}$ odd 1 $b_{1}(1235)$ $1229.5\pm 5.0$ $0\%$ 00
$2^{-+}$ odd 1 $\pi_{2}(1670)$ $1640.57\pm 8.35$ $1.80\%$ $0.39$
$3^{+-}$ odd 1 $b_{3}(2030)$ $1991.55\pm 11$ $1.99\%$ $0.62$
$4^{-+}$ odd 1 $\pi_{4}(2250)$ $2316.73\pm 13$ $2.97\%$ $0.78$
$1^{+-}$ odd 0 $h_{1}(1170)$ $1166\pm 10$ $0\%$ 00
$2^{-+}$ odd 0 $\eta_{2}(1645)$ $1629.93\pm 18$ $0.80\%$ $0.66$
$3^{+-}$ odd 0 $h_{3}(2025)$ $2003.12\pm 25$ $1.08\%$ $1.04$
$4^{-+}$ odd 0 $\eta_{4}(2330)$ $2339.01\pm 30$ $0.47\%$ $1.32$
Figure 1: Regge trajectories (blue lines) for mesons, with masses $m$ expressed in GeV, from the AHW Log model. Red dots are the values obtained from Eq. (23) in the AHW Log model with uncertainties shown in Table 3. The blue lines are obtained by linear regression of these points. Upper left panel: mesons with even ${\cal CP}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${\cal CP}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${\cal CP}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${\cal CP}$, $I=0$ ($h_{1}(1170)$ family).
III.4 AHW model for mesons with odd ${\cal CP}$, $I=0$, and $S\geq 1$
For mesons with odd ${\cal CP}$, $I=0$, and $S\geq 1$, we minimize the function $Q$, Eq. (25), finding $a=0.95\pm 0.06$ for the anomalous dimension, Eq. (22), obtaining
$\Delta_{\rm Log}(S)=(0.95\pm 0.06)\ln(S).$ (33)
The masses obtained in this AHW Log model are presented, with uncertainties, in Table 3, where we used the mass value of the state $h_{1}(1170)$ from PDG as an input, such that $\Lambda_{\rm QCD}=227.04$ MeV. The Regge trajectory obtained by doing linear regression with these masses is given by
$J=(0.73\pm 0.01)M^{2}+0.04\pm 0.05,$ (34)
with $\chi^{2}$/ndf$=3.52/2=1.76.$ This trajectory can be compared with the one obtained from PDG data shown in Table 2 and is presented in the lower right panel of Fig. 1 as the blue line. The red dots are the values of the masses obtained within this AHW Log model with their uncertainties.
@@ IV Linear AHW for mesons
Recently, we have shown Costa-Silva:2023vuu that one can obtain asymptotically linear Regge trajectories for glueballs with even spins $S$ from an AHW model, when considering a total dimension $\Delta_{\rm Total}$ proportional to $\sqrt{S}$. In order to find asymptotically linear Regge trajectories for light unflavored mesons with spins $S\geq 1$, we apply a similar procedure here with a small modification in the power of the spins $S$ as
$\Delta_{\rm Lin}=a_{\rm Lin}(S^{\frac{1}{2}+\epsilon}-1)+4,$ (35)
where in the ideal case, $\epsilon\rightarrow 0$. Note that when $S=1$, which are the lowest spin state for each mesonic family considered here, we recover the canonical dimension $\Delta=4$, compatible with Eq. (11).
Here, the parameter $a_{\rm Lin}$ is obtained exactly as in the AHW Log model, by minimizing $Q$ in Eq. (25). The parameter $\epsilon$ that represents a little deviation of the ideal case $\epsilon=0$ is chosen to obtain the best asymptotic linear Regge trajectory for each meson family. The value of $\epsilon$ is the one that minimizes the total deviation of the masses from the straight line representing the Regge trajectory.
@@
$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm QCD};\qquad\qquad(\nu=\Delta_{\rm Lin}-2),$ (36)
where $\Delta_{\rm Lin}$ is defined in Eq. (35). Since $\Delta_{\rm Lin}=4$ for $S=1$, the values of $\Lambda_{\rm QCD}$ here in this model, for each meson family, are the same as those in the AHW Log model, as well as in the original HW presented in Table 2.
As in the case of logarithmic anomalous dimension, we will analyze the families of light unflavored mesons with even or odd $\cal CP$, isospins $I=0,1$, and spins $S\geq 1$. The Regge trajectories that we obtain within this model are asymptotically linear for each meson family with the appropriate choice of parameters.
IV.1 Linear AHW for mesons with even ${\cal CP}$, $I=1$, and $S\geq 1$
For mesons with even ${\cal CP}$, isospin $I=1$, and spin $S\geq 1$, the parameters in Eq. (35) that minimizes $Q$ are $a_{\rm Lin}=6.87\pm 0.16$ and $\epsilon=0$, so the total dimension reads
$\Delta_{\rm Lin}=(6.87\pm 0.16)\sqrt{S}-2.87\pm 0.16.$ (37)
@@ IV.1 Linear AHW for mesons with even ${\cal CP}$, $I=1$, and $S\geq 1$
with $\chi^{2}$/ndf$=8.68/4=2.17$ and is presented in the upper left panel of Fig. 2 as the blue line. The red dots represent these masses with their uncertainties.
@@
In order to show that the linear AHW model indeed present asymptotic linear Regge trajectories we also plot $J\times M^{2}$ for these mesons with even $\mathcal{CP}$, isospin $I=1$, and $S\geq 1$, for the first 50 states. The trajectory in this case becomes
$J=(0.8479\pm 0.0002)M^{2}+0.60\pm 0.01,$ (39)

### v2

@@ III Anomalous HW model
In an interacting theory, the dimension of an operator is no longer canonical, due to renormalization that introduces anomalous contributions ($\Delta_{\rm Anom}$). So, the total dimension of an operator is given by its canonical plus the anomalous dimensions:
@@
From a well known semi-classical gauge/string duality analysis Gubser:2002tv , it is found that the anomalous dimension increases logarithmically with the spin $S$ of the boundary operators as
@@
where $\lambda$ is the ’t Hooft coupling. Inspired by this result, we consider an anomalous dimension given by
$\Delta_{\rm Log}=a_{\rm Log}\ln(S),$ (20)
where $a_{\rm Log}$ is a constant that will be determined by best fit for each meson family.
@@
$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm HW};\qquad\qquad(\nu=S+1+\Delta_{\rm Log}),$ (21)
where we used the fact that $\Delta_{\rm can.}=3+S$ for these mesonic states and that the order of the corresponding Bessel function is $\nu=\Delta_{\rm Total}-2$. The value of $\Lambda_{\rm HW}$ is determined for each meson familiy using as inputs the masses of the states with $S=1$ from PDG, namely $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$. By construction, these values of $\Lambda_{\rm HW}$ are the same as those ones used in the usual HW model, listed in Table 2, since Eq. (20) implies a vanishing anomalous dimension $\Delta_{\rm Log}$ for $S=1$.
In order to obtain the best numerical fit, let us define the fractional deviation:
$\delta_{i}=\frac{|m_{i}-M_{i}|}{M_{i}},$ (22)
@@
$Q=\sqrt{\sum_{i}\delta_{i}^{2}}.$ (23)
So, the best fit method adopted here corresponds to minimization of the deviations from PDG data given by the function $Q$ for each meson family in the AHW model.
$I^{G}(J^{PC})$ Meson AHW Log $\delta_{\rm Log}$ $\Delta_{\rm Log}$
$1^{+}(1^{--})$ $\rho(770)$ $775.26\pm 10.08$ $0\%$ 00
$1^{-}(2^{++})$ $a_{2}(1320)$ $1309.95\pm 22$ $0.53\%$ $1.92$
$1^{+}(3^{--})$ $\rho_{3}(1690)$ $1681.02\pm 31$ $0.46\%$ $3.04$
$1^{-}(4^{++})$ $a_{4}(1970)$ $1988.75\pm 37$ $1.11\%$ $3.84$
$1^{+}(5^{--})$ $\rho_{5}(2350)$ $2262.37\pm 43$ $2.90\%$ $4.46$
$1^{-}(6^{++})$ $a_{6}(2450)$ $2514.58\pm 47$ $2.64\%$ $4.96$
$0^{-}(1^{--})$ $\omega(782)$ $782.66\pm 10.17$ $0\%$ 00
$0^{+}(2^{++})$ $f_{2}(1270)$ $1302.58\pm 22$ $2.12\%$ $1.81$
$0^{-}(3^{--})$ $\omega_{3}(1670)$ $1666.4\pm 30$ $0.04\%$ $2.87$
$0^{+}(4^{++})$ $f_{4}(2050)$ $1969.63\pm 37$ $2.40\%$ $3.62$
$0^{-}(5^{--})$ $\omega_{5}(2250)$ $2240.19\pm 42$ $0.44\%$ $4.20$
$0^{+}(6^{++})$ $f_{6}(2510)$ $2490.23\pm 46$ $1.02\%$ $4.68$
$1^{+}(1^{+-})$ $b_{1}(1235)$ $1229.5\pm 16$ $0\%$ 00
$1^{-}(2^{-+})$ $\pi_{2}(1670)$ $1640.57\pm 27$ $1.80\%$ $0.39$
$1^{+}(3^{+-})$ $b_{3}(2030)$ $1991.55\pm 37$ $1.99\%$ $0.62$
$1^{-}(4^{-+})$ $\pi_{4}(2250)$ $2316.73\pm 44$ $2.97\%$ $0.78$
$0^{-}(1^{+-})$ $h_{1}(1170)$ $1166\pm 15$ $0\%$ 00
$0^{+}(2^{-+})$ $\eta_{2}(1645)$ $1629.93\pm 25$ $0.80\%$ $0.66$
$0^{-}(3^{+-})$ $h_{3}(2025)$ $2003.12\pm 33$ $1.08\%$ $1.04$
$0^{+}(4^{-+})$ $\eta_{4}(2330)$ $2339.01\pm 39$ $0.47\%$ $1.32$
Table 3: Masses in MeV for mesons classified under even or odd $PC$ with isospins $I=0,1$ and spins $S\geq 1$, from the AHW Log model, determined from Eq. (21), and the percentage deviations $\delta_{\rm Log}=\delta_{i}\times 100\%$, with $\delta_{i}$ given by Eq. (22), of the masses from this model compared with PDG data, using minimization with respect to the parameter $Q$, Eq. (23). For each family the mass of the state with $J=S=1$ is taken as an input from PDG. In the last column we show the anomalous dimension $\Delta_{\rm Log}$ for each state determined by Eq. (20).
Further, we want to know by how much these masses are close to the linear Regge trajectory, obtained from the model, for each family. In order to measure this, we compute
$R^{2}=1-\frac{\sum_{i}[m_{i}-\mathcal{M}_{i}(S)]^{2}}{\sum_{i}(m_{i}-\bar{m})^{2}},$ (24)
where $m_{i}$ is the mass obtained from the AHW model, Eq. (21), for each state $i$, while ${\cal M}_{i}(S)$ is the corresponding mass of this state with angular momentum $S$, computed from the equation of the Regge trajectory and $\bar{m}$ is the average of the masses $m_{i}$. In a perfect model one would have $R^{2}=1$, so the closer $R^{2}$ is to one, the best is the match between the output masses of the model and the linear Regge trajectory.
We also consider an uncertainty for the parameter $a_{\rm Log}$, defined in Eq. (20), which implies uncertainties for the masses ${m_{i}}$, obtained by error propagation, according to the formula:
@@
where $\delta m_{0}$ and $\delta\chi_{\nu_{i},1}$ are the uncertainties in the mass of the state with $S=1$ and in the zero of Bessel function that is related to the mass $m_{i}$. The uncertainty $\delta m_{0}$ of the state with $S=1$ for each meson family in our model will be the computed as follows: we consider $\delta m_{0}$ in the family $\rho(770)$ as the arithmetic mean of all uncertainties in the PDG data for this family, which represents 1.3% of the mass of $\rho(770)$. In order to have a uniform criterium, we will take the uncertainties $\delta m_{0}$ of each family to be 1.3% of their $S=1$ state mass.
$I^{G}(J^{PC})$ Meson $a_{\rm Log}$ AHW (Log) Regge Trajectory (GeV) $R^{2}$
$1^{+}(1^{--})$ $\rho(770)$ $2.77\pm 0.12$ $J=(0.88\pm 0.01)M^{2}+0.50\pm 0.03$ 0.9998
$0^{-}(1^{--})$ $\omega(782)$ $2.61\pm 0.09$ $J=(0.90\pm 0.01)M^{2}+0.48\pm 0.03$ 0.9997
$1^{+}(1^{+-})$ $b_{1}(1235)$ $0.56\pm 0.03$ $J=(0.78\pm 0.02)M^{2}-0.13\pm 0.08$ 0.9985
$0^{-}(1^{+-})$ $h_{1}(1170)$ $0.95\pm 0.06$ $J=(0.73\pm 0.01)M^{2}+0.04\pm 0.05$ 0.9993
Table 4: The Regge trajectories in GeV units for each light unflavored meson family (same $PC$ parity and same isospin $I$) from the AHW model, Eqs. (20) and (21), with the corresponding values of $R^{2}$ from Eq. (24). The coefficients $a_{\rm Log}$ presented here come from the minimization of $Q$, Eq. (23).
Now that we have defined our criteria for best fit and how we compute the uncertainties $\delta{m_{i}}$ for each calculated mass $m_{i}$, we are able to tackle the AHW model for light unflavored mesons.
Figure 2: Regge trajectories ($J\times M^{2}$) for mesons, with masses $M$ expressed in GeV, from the AHW Log model (blue lines) using the minimization of $Q$, Eq. (23), and from PDG data (orange lines). Red dots are the values obtained from Eq. (21), while black dots from PDG data, both shown with uncertainties in Tables 3 and 1, respectively. The lines are obtained by linear regression of these points. Upper left panel: mesons with even ${PC}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${PC}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${PC}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${PC}$, $I=0$ ($h_{1}(1170)$ family).
We start applying the AHW model for mesons with even ${PC}$, isospin $I=1$, and spins $S\geq 1$, taking the mass of the $\rho(770)$ state from PDG as an input, which fixes $\Lambda_{\rm HW}=150.96$ MeV for this family. We minimize the parameter $Q$, given by Eq. (23), with respect to PDG masses from Table 1, finding the logarithm anomalous dimensions, Eq. (20), to be
@@
The masses with their uncertainties obtained within this model, from Eq. (21), with these anomalous dimensions, are presented in Table 3, together with percentage deviations in comparison with PDG data. The value of $\Delta_{\rm Log}(S)$ of the above equation for each spin $S$ is also shown in the last column of this Table. The Regge trajectory obtained by doing linear regression with the obtained masses from this AHW Log model is given by
$J=(0.88\pm 0.01)M^{2}+0.50\pm 0.03.$ (27)
@@
This trajectory is presented in the upper left panel of Fig. 2 as the blue line, together with the one from PDG (orange line). The red dots are the values in the $J\times M^{2}$ plane of the mesons $\rho(770)$, $a_{2}(1320)$, $\rho_{3}(1690)$, $a_{4}(1970)$, $\rho_{5}(2350)$, and $a_{6}(2450)$ in this AHW Log model, with their uncertainties, while the black dots represent PDG data from Table 1.
[…]
For completeness and comparison, we also present in the Appendix VI a more standard minimization of $\chi^{2}$, defined in Eq. (33). As one can check, the results of both minimizations are very close to each other (actually, they coincide for the $h_{1}(1170)$ family, within the precision we are working here).
@@ IV Linear AHW for mesons
Recently, we have shown Costa-Silva:2023vuu that one can obtain asymptotically linear Regge trajectories for glueballs with even spins $S$ from an AHW model, when considering a total dimension, $\Delta_{\rm Total}$, proportional to $\sqrt{S}$. By asymptotically linear Regge trajectories, we mean functional relations between $J$ and $M^{2}$ which are not exactly linear but can approximate very well the linear behavior with the desired precision.
In order to find asymptotically linear Regge trajectories for light unflavored mesons with spins $S\geq 1$, we apply a similar procedure here with a small modification in the power of the spins $S$ as
$\Delta_{\rm LAHW}=a_{\rm LAHW}(S^{\frac{1}{2}+\epsilon}-1)+4,$ (28)
where in the ideal case, $\epsilon\rightarrow 0$. Note that when $S=1$, which is the lowest spin state for each mesonic family considered here, we recover the canonical dimension $\Delta=4$, compatible with Eq. (9).
$I^{G}(J^{PC})$ Meson Linear AHW $\delta_{\rm LAHW}$ $\Delta_{\rm LAHW}$
$1^{+}(1^{--})$ $\rho(770)$ $775.26\pm 10.08$ $0\%$ 00
$1^{-}(2^{++})$ $a_{2}(1320)$ $1296.75\pm 21$ $1.53\%$ $1.85$
$1^{+}(3^{--})$ $\rho_{3}(1690)$ $1678.61\pm 30$ $0.60\%$ $3.03$
$1^{-}(4^{++})$ $a_{4}(1970)$ $1993.84\pm 38$ $1.36\%$ $3.87$
$1^{+}(5^{--})$ $\rho_{5}(2350)$ $2268.03\pm 44$ $2.66\%$ $4.49$
$1^{-}(6^{++})$ $a_{6}(2450)$ $2513.71\pm 51$ $2.60\%$ $4.96$
$0^{-}(1^{--})$ $\omega(782)$ $782.66\pm 10.17$ $0\%$ 00
$0^{+}(2^{++})$ $f_{2}(1270)$ $1294.28\pm 20$ $1.47\%$ $1.76$
$0^{-}(3^{--})$ $\omega_{3}(1670)$ $1669.09\pm 29$ $0.13\%$ $2.88$
$0^{+}(4^{++})$ $f_{4}(2050)$ $1978.53\pm 37$ $1.96\%$ $3.67$
$0^{-}(5^{--})$ $\omega_{5}(2250)$ $2247.67\pm 44$ $0.10\%$ $4.24$
$0^{+}(6^{++})$ $f_{6}(2510)$ $2488.81\pm 50$ $0.97\%$ $4.67$
$1^{+}(1^{+-})$ $b_{1}(1235)$ $1229.5\pm 16$ $0\%$ 00
$1^{-}(2^{-+})$ $\pi_{2}(1670)$ $1667.30\pm 26$ $0.20\%$ $0.48$
$1^{+}(3^{+-})$ $b_{3}(2030)$ $2000.60\pm 37$ $1.55\%$ $0.65$
$1^{-}(4^{-+})$ $\pi_{4}(2250)$ $2281.03\pm 46$ $1.38\%$ $0.65$
$0^{-}(1^{+-})$ $h_{1}(1170)$ $1166\pm 15$ $0\%$ 00
$0^{+}(2^{-+})$ $\eta_{2}(1645)$ $1646.87\pm 26$ $1.85\%$ $0.72$
$0^{-}(3^{+-})$ $h_{3}(2025)$ $2009.96\pm 37$ $0.74\%$ $1.07$
$0^{+}(4^{-+})$ $\eta_{4}(2330)$ $2314.19\pm 47$ $0.59\%$ $1.22$
Table 5: Masses in MeV for mesons classified under even or odd $PC$ with isospins $I=0,1$ and spins $S\geq 1$, from the Linear AHW model, determined from Eq. (29), and the percentage deviations $\delta_{\rm LAHW}=\delta_{i}\times 100\%$, with $\delta_{i}$ given by Eq. (22), of the masses from this model compared with PDG data, using minimization with respect to the parameter $Q$, Eq. (23). For each family the mass of the state with $J=S=1$ is taken as an input. In the last column we show the total dimension $\Delta_{\rm LAHW}$ for each state determined by Eq. (28).
$I^{G}(J^{PC})$ Meson $a_{\rm LAHW}$ $\epsilon$ Linear AHW Regge Trajectory (GeV) $n$ $R^{2}$
$1^{+}(1^{--})$ $\rho(770)$ $6.87\pm 0.16$ $0.00\pm 0.005$ $J=(0.87\pm 0.01)M^{2}+0.51\pm 0.02$ 6 0.999793
$J=(0.8479\pm 0.0002)M^{2}+0.60\pm 0.01$ 50 0.999997
$0^{-}(1^{--})$ $\omega(782)$ $6.67\pm 0.13$ $0.00\pm 0.005$ $J=(0.89\pm 0.01)M^{2}+0.49\pm 0.02$ 6 0.999845
$J=(0.8766\pm 0.0003)M^{2}+0.52\pm 0.01$ 50 0.999995
$1^{+}(1^{+-})$ $b_{1}(1235)$ $3.19\pm 0.03$ $0.05\pm 0.005$ $J=(0.81\pm 0.01)M^{2}-0.24\pm 0.03$ 4 0.999846
$J=(0.8549\pm 0.0003)M^{2}-0.42\pm 0.01$ 50 0.999993
$0^{-}(1^{+-})$ $h_{1}(1170)$ $3.79\pm 0.08$ $0.04\pm 0.005$ $J=(0.75\pm 0.03)M^{2}-0.03\pm 0.01$ 4 0.999960
$J=(0.7667\pm 0.0003)M^{2}-0.006\pm 0.01$ 50 0.999994
Table 6: The Regge trajectories in GeV units for each light unflavored meson family (same $PC$ parity and same isospin $I$) from the Linear AHW model, Eq. (29). For each meson family we show two Regge trajectories corresponding to the linear regression of two sets of points with dimension $n$. These trajectories are plot in Figures 3 and 4. We also show the corresponding values of $R^{2}$ from Eq. (24). The coefficients $a_{\rm LAHW}$ and $\epsilon$ presented here come from the minimization of $Q$, Eq. (23).
Here, the parameter $a_{\rm LAHW}$ is obtained exactly as in the AHW Log model, by minimizing $Q$ in Eq. (23). The parameter $\epsilon$ that represents a little deviation of the ideal case $\epsilon=0$ is chosen to obtain the best asymptotically linear Regge trajectory for each meson family. The value of $\epsilon$ is the one that minimizes the total deviation of the masses from the straight line representing the Regge trajectory.
@@
$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm HW};\qquad\qquad(\nu=\Delta_{\rm LAHW}-2),$ (29)
where $\Delta_{\rm LAHW}$ is defined in Eq. (28). Since $\Delta_{\rm LAHW}=4$ for $S=1$, the values of $\Lambda_{\rm HW}$ here in this model, for each meson family, are the same as those in the AHW Log model, as well as in the original HW presented in Table 2.
As in the case of logarithmic anomalous dimension, we will analyze the families of light unflavored mesons with even or odd $PC$, isospins $I=0,1$, and spins $S\geq 1$. The Regge trajectories that we obtain within this model are asymptotically linear for each meson family with the appropriate choice of parameters.
Figure 3: Regge trajectories ($J\times M^{2}$) for mesons, with masses $M$ expressed in GeV, from the Linear AHW model (blue lines) with total dimension given by Eq. (28), using the minimization of $Q$, Eq. (23), and from PDG data (orange lines). Red dots are the values obtained from Eq. (29), while black dots come from PDG, both with uncertainties. The trajectories are obtained by linear regression of these points, for each meson family. Upper left panel: mesons with even $PC$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${PC}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${PC}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${PC}$, $I=0$ ($h_{1}(1170)$ family).
Figure 4: Regge trajectories ($J\times M^{2}$) for mesons, with masses $M$ expressed in GeV, from the linear AHW model, with total dimension given by Eq. (28), using the minimization of $Q$, Eq. (23). Red dots are the values obtained from Eq. (29) with uncertainties and the blue lines are the Regge trajectories, obtained by linear regression of these 50 points, for each meson family. Upper left panel: mesons with even ${PC}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${PC}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${PC}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${PC}$, $I=0$ ($h_{1}(1170)$ family).
As in the previous section, we start with the $\rho(770)$ family, i.e., for mesons with even ${PC}$, isospin $I=1$, and spin $S\geq 1$, the parameters in Eq. (28) that minimize $Q$ are $a_{\rm LAHW}=6.87\pm 0.16$ and $\epsilon=0.00\pm 0.005$, so the total dimension reads
$\Delta_{\rm LAHW}=(6.87\pm 0.16)\sqrt{S}-2.87\pm 0.16.$ (30)
@@
and is presented in the upper left panel of Fig. 3 as the blue line, together with the PDG data represented by black dots with Regge trajectory given by the orange line. The red dots represent the masses of the LAHW model with their uncertainties.
@@
In order to show that the linear AHW model indeed presents asymptotically linear Regge trajectories we also plot $J\times M^{2}$ for these mesons with even $PC$, isospin $I=1$, and $S\geq 1$, for the first 50 states with masses calculated from this model. The trajectory in this case becomes
$J=(0.8479\pm 0.0002)M^{2}+0.60\pm 0.01.$ (32)